# HAODiff: Human-Aware One-Step Diffusion via Dual-Prompt Guidance

### v1

@@ 5 Limitation and Conclusion
LQ SUPIR yu2024supir HAODiff (ours)
@@
Limitation. Images with heavy global degradation and strong local motion blur, particularly in regions with fast-moving limbs, remain challenging for all methods, including ours. A few samples of this type appear in the MPII-Test. Although HAODiff outperforms all baselines (many of them fail completely, as shown in Fig. 8 and supplementary material), our results remain imperfect. Limb poses may appear unnatural, and overlapping objects may be misrestored. These limitations prompt us to reconsider the boundary between pure restoration and content generation. Moreover, similar to common CFG strategies in image restoration, our method employs UNet to compute the latent vector twice. These two passes thus incur a modest efficiency penalty compared to OSEDiff wu2024osediff and OSDHuman gong2025osdhuman (see Tab. 1). In future work, we will develop more efficient CFG strategies and explore approaches to handle such extreme real-world cases, recover finer details.

### v2

@@ 5 Limitation and Conclusion
LQ SUPIR Yu et al. (2024) HAODiff (ours)
@@
Limitation. Images with heavy global degradation and strong local motion blur, particularly in regions with fast-moving limbs, remain challenging for all methods, including ours. A few samples of this type appear in the MPII‑Test. Although HAODiff outperforms all baselines (many of them fail completely, as shown in Fig. 8 and supplementary material), our results remain imperfect. Limb poses may appear unnatural, and overlapping objects may be misrestored. These limitations prompt us to reconsider the boundary between pure restoration and content generation. Moreover, similar to common CFG strategies in image restoration, our method employs UNet to compute the latent vector twice. These two passes thus incur a modest efficiency penalty compared to OSEDiff Wu et al. (2024a) and OSDHuman Gong et al. (2025) (see Tab. 1). In future work, we will develop more efficient CFG strategies and explore approaches to handle such extreme real‑world cases, recover finer details.
@@
Acknowledgments
This work was supported by Shanghai Municipal Science and Technology Major Project (2021SHZDZX0102), the Fundamental Research Funds for the Central Universities, the Special Project on Technological Innovation Application for the 15th National Games and the National Paralympic Games under Grant 2025B01W0005.